# Causal Graph Neural Networks for Healthcare

### v3

@@ 3.5 Ensuring causal fairness in clinical algorithms
The RaVSNet framework demonstrates the clinical impact of causal fairness correction through explicit modelling of prescription bias [71]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independently of therapeutic benefit. The system generates causal effect matrices, adjusting the probabilities of recommendations to reflect genuine therapeutic value rather than the tradition of prescribing. On MIMIC-III and MIMIC-IV datasets, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively, representing 3–7% improvements over baseline methods, while the causal effect inference module specifically addresses confounding from historical prescribing patterns that may disadvantage underserved populations [71].
@@ 4.2 Prognosis: predicting what interventions prevent
Prognostic models traditionally predict outcomes based on baseline characteristics, yet clinical utility requires understanding which interventions might prevent adverse events. The Knowledge-aware Multi-label Network addresses a critical limitation in complication prediction by explicitly modelling causal relationships between complications rather than treating them as independent events [78]. The framework constructs knowledge graphs depicting causal relationships such as ”lung infection causes pleural effusion,” which, in turn, increases the risk of respiratory failure. This understanding of the causal chain enables clinicians to implement preventive measures targeting the root causes rather than managing complications reactively.
In a cohort of 593 patients with lung cancer surgery, KAMLN achieved micro-AUC of 0.664 for multi-label complication prediction by explicitly modelling causal dependencies between complications [78]. The knowledge-aware architecture enables earlier identification of causal chains, supporting prophylactic interventions before complications cascade. The interpretability of the model, showing not only risk, but also causal pathways, helped clinicians understand and trust the predictions, which proved crucial for clinical adoption. The explicit modelling of causal dependencies between complications has important equity implications beyond predictive performance. Patients in underserved communities, who often have more advanced disease states, benefit disproportionately from early identification of causal chains.
@@ 4.3 Treatment selection: mechanism-based therapy
Drug recommendation in the era of polypharmacy requires understanding not just individual drug effects but also complex interactions and patient-specific factors. The Conditional Causal Representation Learner addresses spurious correlations that plague traditional recommendation systems [68]. For example, proton pump inhibitors might appear beneficial for cardiovascular outcomes not due to their therapeutic effect but because they are prescribed to patients who receive aspirin for cardioprotection. CRec disentangles these confounded relationships by separating drug molecular representations into causal therapeutic components and spurious prescription patterns. Through causal intervention to eliminate spurious effects, recommendations become genuinely mechanism-focused. In a study of 47,326 patients with multiple chronic conditions, CRec identified drug combinations that traditional methods failed to identify because they were rarely co-prescribed despite mechanistic synergy [68]. One notable discovery involved the combination of a diabetes medication with an Alzheimer’s drug based on shared metabolic pathways, a combination now in clinical trials showing promising results for both conditions.
@@
RaVSNet improves the precision of the recommendation by explicitly modelling and correcting prescription bias through Causal Effect Inference tasks [71]. Using causal discovery through Bayesian networks and GIES algorithms, it quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independent of therapeutic benefit. The system generates causal effect matrices that adjust recommendation probabilities to reflect genuine therapeutic value rather than prescribing tradition. On MIMIC-III, RaVSNet achieved PRAUC of 0.814 and Jaccard of 0.573, improvements of 4–7% over state-of-the-art baselines, while the causal framework provides a principled mechanism for identifying and correcting prescription biases that may systematically disadvantage specific patient populations [71].
@@
For drug discovery, PDGrapher uses protein-protein interaction networks as approximations of causal graphs, operating under the assumption that there are no unobserved confounders to predict combinatorial therapeutic targets [50]. This causally-inspired approach outperforms traditional methods in identifying effective perturbagens across nine cell lines with chemical perturbations. The identified causal targets provide more than drug efficacy predictions. In cancer applications, PDGrapher revealed synthetic lethality relationships in which targeting two proteins simultaneously produces therapeutic benefit despite individual targets showing minimal effect. These mechanistic insights enable the rational design of combination therapies, accelerating the path from target identification to clinical translation. The clinical workflow transforms from ”drug prescribed for patients like you” to ”drug targeting your causal mechanisms,” representing a fundamental shift toward mechanism-based precision medicine.
@@ 4.4 Continuous monitoring: real-time causal inference
In spatiotemporal time series imputation, critical for handling missing data in clinical monitoring, the Casper framework addresses a fundamental challenge of unknown confounders that create non-causal shortcut backdoor paths [80]. For example, the time of day might influence both the monitoring frequency and physiological parameters, creating spurious correlations. Casper applies frontdoor adjustment to eliminate these confounding paths while preserving genuine causal relationships. The Spatiotemporal Causal Attention module of the framework enforces sparse, interpretable causality among embeddings, with theoretical guarantees of convergence to gradient-based explanations [80]. In multi-site sensor network data, Casper achieved substantial improvements in imputation quality compared to traditional methods, with lower MAE and MSE scores across all evaluated datasets while maintaining reduced variance, demonstrating that enforcing causal attention can improve robustness. This improvement translates directly to clinical outcomes through more accurate imputation, thereby improving the prediction of deterioration and reducing unnecessary interventions triggered by imputation errors.
@@
Complication prediction KAMLN [78] Knowledge graph encoding causal dependencies between complications Micro-AUC 0.664 for multi-label prediction; causal chain modelling enables targeted prophylaxis
Drug recommendation CRec [68], RaVSNet [71] Disentanglement of therapeutic effects from prescription bias Identified mechanistically synergistic combinations; 4–7% improvement in recommendation accuracy via causal bias correction
Combination therapy PDGrapher[50], RC-Explainer[60] Causal inference on protein-protein interaction networks Synthetic lethality identification; drug synergy prediction from observational data
@@
ICU data imputation Casper [80] Frontdoor adjustment for spatiotemporal confounding Reduced MAE/MSE with lower variance across multi-site sensor data; causal validity under intervention
@@ 5.1 Patient-specific mechanistic simulation
Current methodological advances establish foundations for realising this vision. Interventional GNNs such as iVGAE and RC-Explainer enable simulation of treatments never observed in training data by implementing do-calculus through differentiable neural operations [45, 60]. Counterfactual generators, including CLEAR and CXGNN, explore alternative scenarios while maintaining biological plausibility, computing patient-specific predictions through the abduction-action-prediction framework [54, 66]. Multi-modal integration frameworks such as MOGONET and MoCaGCN combine diverse data modalities through shared causal structure, learning representations that capture biological mechanisms spanning molecular to phenotypic scales [77, 82]. Fairness constraints ensure equitable simulation by identifying and mitigating discriminatory causal pathways, preventing digital twins from perpetuating biases embedded in training data [13]. The technical infrastructure exists; clinical validation and regulatory frameworks remain the primary barriers to deployment.
@@ 6.4 From methods to clinical practice
Box 3: Example of multi-modal validation framework The amyloid-$\beta$ hypothesis in Alzheimer’s disease Consider a causal GNN identifying a directed edge from amyloid-$\beta$ deposition to cognitive decline in Alzheimer’s disease, a relationship with validation complexities. Evidence triangulation: 1. Biological plausibility ($S_{\text{bio}}=+1$): Strong mechanistic support through the amyloid cascade hypothesis and decades of foundational neuroscience research. 2. Replication ($S_{\text{rep}}=+1$): Association replicates consistently across major cohorts, including ADNI, AIBL, and European EADC datasets with stable effect sizes. 3. Natural experiments ($S_{\text{quasi}}=-1$): Mendelian randomization studies using genetic variants affecting amyloid production demonstrate no causal effect on cognitive decline [103]. 4. Prospective validation ($S_{\text{prosp}}=-1$): Multiple Phase III clinical trials of amyloid-reducing therapies show minimal cognitive benefit despite substantial plaque reduction [104]. 5. Sensitivity analysis ($S_{\text{sens}}=0$): Moderate robustness to unmeasured confounding (E-value = 1.8) but substantial uncertainty regarding alternative causal structures. Aggregate evidence score: $\text{AES}=0.20(+1)+0.25(+1)+0.30(-1)+0.15(-1)+0.10(0)=0.00$ (4) Interpretation: Despite strong observational associations and biological plausibility, the null aggregate score reflects fundamentally contradictory evidence from high-quality natural experiments and prospective trials [97, 105]. This suggests the observational causal model may be confounded, with amyloid deposition potentially representing a biomarker of disease progression rather than a causal driver, or causality involving more complex dynamics, including threshold effects or critical developmental periods that the learned graph structure fails to capture. Clinical implication: Predictions based on this causal edge should be treated with appropriate caution, and amyloid reduction alone may not constitute an effective therapeutic strategy. This exemplifies how systematic evidence triangulation prevents overconfident causal claims despite compelling observational evidence, ultimately protecting patients from ineffective interventions whilst guiding research toward more productive therapeutic targets.
@@ 7 Conclusion
Applications spanning psychiatric subtyping, cancer molecular characterisation, physiological monitoring, and drug discovery validate this approach across diverse clinical domains. The CI-GNN framework maintained stable diagnostic performance across institutions where traditional methods varied dramatically [52]. Multi-omics causal integration identified 89 driver genes from thousands of correlational features, accelerating precision oncology [82]. Causal feature selection transformed continuous monitoring from high-dimensional noise to mechanistically interpretable signals [51]. Treatment recommendation systems that correct for prescription bias achieved substantial outcome improvements in historically underserved populations [71]. The methodological foundations for patient-specific Causal Digital Twins now exist, enabling in silico clinical experimentation before therapeutic intervention, with synergistic integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation, positioning the field at the forefront of multimodal artificial intelligence [31, 87].
@@
Healthcare artificial intelligence is at an inflexion point. The question is no longer whether machine learning can achieve impressive metrics in retrospective datasets but whether systems can survive deployment across diverse clinical contexts while maintaining effectiveness, equity, and interpretability. Causal graph neural networks represent the only architectural paradigm that simultaneously addresses robustness through invariant learning, fairness through path-specific interventional constraints, and mechanistic understanding through explicit causal structure [28, 67, 12]. Success requires coordinated action on multiple fronts. Computational infrastructure investments must enable hierarchical causal discovery exploiting biological modularity, amortised inference achieving clinical response times, and neuromorphic hardware acceleration for graph operations at scale [81]. Validation standards require community consensus on evidentiary requirements for causal claims from observational data, analogous to established frameworks for randomised trials but adapted to healthcare’s unique constraints. Regulatory pathways must evolve to evaluate systems that make interventional predictions and counterfactual reasoning, extending beyond traditional accuracy-based assessment to mechanistic validity [33]. Clinical trials demonstrating impact on patient outcomes rather than surrogate metrics prove essential to justify substantial implementation costs. Training the next generation of clinician-scientists fluent in both causal inference theory and clinical medicine will determine whether these methods reach their potential.

### v4

@@ 3.5 Ensuring causal fairness in clinical algorithms
The RaVSNet framework demonstrates the clinical necessity of causal bias correction by explicitly modelling the discrepancy between historical prescription habits and therapeutic efficacy [66]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how institutional protocols and co-occurrence patterns influence medication choices independently of patient-specific benefit. The system generates causal effect matrices to adjust recommendation probabilities, filtering out spurious associations driven by data availability rather than medical necessity. On MIMIC-III and MIMIC-IV, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively (3–7% improvement), while its causal inference module specifically mitigates the structural confounding in electronic health records, an important step for preventing the propagation of historical inequities that often manifest as data sparsity or noise in underserved populations [66].
@@ 4.2 Prognosis: predicting what interventions prevent
The Knowledge-aware Multi-label Network (KAMLN) addresses a significant limitation in postoperative complication prediction by explicitly modelling dependencies between complications rather than treating them as independent events [73]. The framework utilizes a heterogeneous knowledge graph depicting clinical relationships such as ”lung infection causes pleural effusion,” which in turn is linked to atelectasis [73]. In a cohort of 593 lung cancer surgery patients, KAMLN achieved a micro-AUC of 0.664, demonstrating that integrating prior medical knowledge can stabilize predictions for rare complications—such as pulmonary embolism and deep vein thrombosis—where traditional single-label models often fail [73]. The model’s use of SHAP analysis and a five-layer graph neural network architecture provides a transparent approach to understanding the clinical features that drive multi-label risk, identifying factors like operation duration and the number of dissected lymph nodes as critical predictors across multiple complication groups [73]. While further external validation is required before direct clinical application, the explicit modelling of these causal chains facilitates more accurate and fine-grained risk assessment for surgical patients [73].
@@ 4.3 Treatment selection: mechanism-based therapy
Drug recommendation in the era of polypharmacy requires understanding not just individual drug effects but also complex interactions and patient-specific factors. The Conditional Causal Representation Learner tries to tackle spurious correlations that plague traditional recommendation systems [63]. For example, proton pump inhibitors might appear beneficial for cardiovascular outcomes not due to their therapeutic effect but because they are prescribed to patients who receive aspirin for cardioprotection. CRec disentangles these confounded relationships by separating drug molecular representations into causal therapeutic components and spurious prescription patterns. Through causal intervention to eliminate spurious effects, recommendations become genuinely mechanism-focused. In a study of 47,326 patients with multiple chronic conditions, CRec identified drug combinations that traditional methods failed to identify because they were rarely co-prescribed despite mechanistic synergy [63]. One notable discovery involved the combination of a diabetes medication with an Alzheimer’s drug based on shared metabolic pathways, a combination now in clinical trials showing promising results for both conditions.
@@
RaVSNet improves the precision of the recommendation by explicitly modelling and correcting data distribution bias through Causal Effect Inference tasks [66]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how clinical factors (diagnoses, symptoms) influence medication choices independent of simple co-occurrence patterns. The system generates causal effect matrices that adjust recommendation probabilities to reflect genuine therapeutic associations rather than spurious statistical correlations. On MIMIC-III and MIMIC-IV datasets, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively, improvements over state-of-the-art baselines, while the causal framework provides a principled mechanism for identifying and correcting confounding biases in retrospective electronic health records [66].
@@
For drug discovery, PDGrapher uses protein-protein interaction networks as approximations of causal graphs, operating under the assumption that there are no unobserved confounders to predict combinatorial therapeutic targets [47]. This causally-inspired approach outperforms traditional methods in identifying effective perturbagens across nine cell lines with chemical perturbations. The identified causal targets provide more than simple efficacy predictions; by solving the inverse problem of perturbagen discovery, the model reveals combinatorial dependencies where targeting multiple specific proteins is required to shift a unique disease state to a healthy profile. These mechanistic insights could allow the rational design of combination therapies that exploit biological structure rather than trial-and-error. Consequently, the clinical workflow transforms from ”drug prescribed for patients like you” to ”drug targeting your causal mechanisms,” representing a fundamental shift toward mechanism-based precision medicine.
@@ 4.4 Continuous monitoring: real-time causal inference
In high-stakes clinical monitoring, missing data is often the reality which spatio-temporal modelling has to overcome. To this end, the Casper framework can provide a proof-of-concept to addresses the challenge of unknown confounders that create non-causal shortcut backdoor paths [75]. While evaluated on sensor networks, the framework’s use of frontdoor adjustment offers a methodology to eliminate spurious correlations, such as those created by hospital environmental factors influencing both monitoring frequency and patient physiological parameters, while preserving genuine causal relationships [75]. The Spatiotemporal Causal Attention module of the framework enforces sparse, interpretable causality among embeddings, with theoretical guarantees of convergence to gradient-based explanations [75]. By demonstrating substantial improvements in imputation quality and reduced variance in complex multi-site datasets, Casper provides a robust foundation for more accurate clinical data reconstruction [75]. This enhanced accuracy potentially improves the prediction of patient deterioration and reduces the likelihood of unnecessary interventions triggered by imputation errors, effectively translating causal discovery into improved clinical outcomes.
@@
Complication prediction KAMLN [73] Heterogeneous knowledge graph encoding clinical dependencies between complications Micro-AUC 0.664 for multi-label prediction; SHAP analysis identifies lymph node dissection as a significant risk factor
Drug recommendation CRec [63], RaVSNet [66] Disentanglement of therapeutic effects from prescription bias Identified mechanistically synergistic combinations (CRec); 4–7% accuracy improvement over baselines with reduced DDI rates (RaVSNet)
Combination therapy PDGrapher [47] Inverse causal inference on protein-protein interaction networks Direct prediction of combinatorial perturbagens for phenotype reversal; 13.3% improvement in identifying therapeutic targets from interventional data
@@
ICU data imputation Casper [75] Frontdoor adjustment to block unknown spatiotemporal confounders Significantly reduced MAE/MSE; methodological template for handling non-causal shortcuts in continuous monitoring
@@ 5.1 Patient-specific mechanistic simulation
Current methodological advances establish foundations for realising this vision. Interventional GNNs such as iVGAE enable simulation of treatments never observed in training data by implementing do-calculus through differentiable neural operations [42]. Counterfactual generators, including CLEAR and CXGNN, explore alternative scenarios while maintaining biological plausibility, computing patient-specific predictions through the abduction-action-prediction framework [51, 60]. Multi-modal integration frameworks such as MOGONET and MoCaGCN combine diverse data modalities through shared causal structure, learning representations that capture biological mechanisms spanning molecular to phenotypic scales [72, 77]. Fairness constraints ensure equitable simulation by identifying and mitigating discriminatory causal pathways, preventing digital twins from perpetuating biases embedded in training data [11].
@@ 6.4 From methods to clinical practice
Box 3: Example of multi-modal validation framework The amyloid-$\beta$ hypothesis in Alzheimer’s disease Consider a causal GNN identifying a directed edge from amyloid-$\beta$ deposition to cognitive decline in Alzheimer’s disease, a relationship with validation complexities. Evidence triangulation: 1. Biological plausibility ($S_{\text{bio}}=+1$): Strong mechanistic support through the amyloid cascade hypothesis and decades of foundational neuroscience research. 2. Replication ($S_{\text{rep}}=+1$): Association replicates consistently across major cohorts, including ADNI, AIBL, and European EADC datasets with stable effect sizes. 3. Natural experiments ($S_{\text{quasi}}=-1$): Mendelian randomization studies using genetic variants affecting amyloid production demonstrate no causal effect on cognitive decline. 4. Prospective validation ($S_{\text{prosp}}=-1$): Multiple Phase III clinical trials of amyloid-reducing therapies show minimal cognitive benefit despite substantial plaque reduction. 5. Sensitivity analysis ($S_{\text{sens}}=0$): Moderate robustness to unmeasured confounding (E-value = 1.8) but substantial uncertainty regarding alternative causal structures. Aggregate evidence score: $\text{AES}=0.20(+1)+0.25(+1)+0.30(-1)+0.15(-1)+0.10(0)=0.00$ (4) Interpretation: Despite strong observational associations and biological plausibility, the null aggregate score could reflect contradictory evidence from high-quality natural experiments and prospective trials [92, 98]. This suggests the observational causal model may be confounded, with amyloid deposition potentially representing a biomarker of disease progression rather than a causal driver, or causality involving more complex dynamics, including threshold effects or critical developmental periods that the learned graph structure fails to capture. Clinical implication: Predictions based on this causal edge should be treated with appropriate caution, and amyloid reduction alone may not constitute an effective therapeutic strategy. This exemplifies how systematic evidence triangulation prevents overconfident causal claims despite compelling observational evidence, ultimately protecting patients from ineffective interventions whilst guiding research toward more productive therapeutic targets.
@@ 7 Conclusion
Applications spanning psychiatric subtyping, cancer molecular characterisation, physiological monitoring, and drug discovery validate this approach across diverse clinical domains. The CI-GNN framework maintained stable diagnostic performance across institutions where traditional methods varied dramatically [50]. Multi-omics causal integration identified 89 driver genes from thousands of correlational features, accelerating precision oncology [77]. Causal feature selection transformed continuous monitoring from high-dimensional noise to mechanistically interpretable signals [48]. The methodological foundations for patient-specific Causal Digital Twins now exist, enabling in silico clinical experimentation before therapeutic intervention, with synergistic integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation, positioning the field at the forefront of multimodal artificial intelligence [29, 82].
@@
Healthcare artificial intelligence is at an inflexion point. The question is no longer whether machine learning can achieve impressive metrics in retrospective datasets but whether systems can survive deployment across diverse clinical contexts while maintaining effectiveness, equity, and interpretability. Causal graph neural networks represent the only architectural paradigm that simultaneously addresses robustness through invariant learning, fairness through path-specific interventional constraints, and mechanistic understanding through explicit causal structure [26, 62, 10]. Success requires coordinated action on multiple fronts. Computational infrastructure investments must enable hierarchical causal discovery exploiting biological modularity, amortised inference achieving clinical response times, and neuromorphic hardware acceleration for graph operations at scale [76]. Validation standards require community consensus on evidentiary requirements for causal claims from observational data, analogous to established frameworks for randomised trials but adapted to healthcare’s unique constraints. Regulatory pathways must evolve to evaluate systems that make interventional predictions and counterfactual reasoning, extending beyond traditional accuracy-based assessment to mechanistic validity. Clinical trials demonstrating impact on patient outcomes rather than surrogate metrics prove essential to justify substantial implementation costs. Training the next generation of clinician-scientists fluent in both causal inference theory and clinical medicine will determine whether these methods reach their potential.